### v1

# Ahmet Faruk Saz,  Siheng Xiong,  and Faramarz Fekri

Abstract
This work presents an analysis of semantic communication in the context of First-Order Logic (FOL)-based deduction. Specifically, the receiver holds a set of hypotheses about the State of the World (SotW), while the transmitter has incomplete evidence about the true SotW but lacks access to the ground truth. The transmitter aims to communicate limited information to help the receiver identify the hypothesis most consistent with true SotW. We formulate the objective as approximating the posterior distribution at the transmitter to the receiver. Using Stirling’s approximation, this reduces to a constrained, finite-horizon resource allocation problem. Applying the Karush-Kuhn-Tucker conditions yields a truncated water-filling solution. Despite the problem’s non-convexity, symmetry and permutation invariance ensure global optimality. Based on this, we design message selection strategies, both for single- and multi- round communication, and model the receiver’s inference as an $m$-ary Bayesian hypothesis testing problem. Under the Maximum A Posteriori (MAP) rule, our communication strategy achieves optimal performance within budget constraints. We further analyze convergence rates and validate the theoretical findings through experiments, demonstrating reduced error over random selection and prior methods.
Index Terms:
First-order logic, semantic communication, water-filling, hypothesis testing, deduction
I Introduction
Semantic communication redefines data transmission by prioritizing meaning over raw data, offering critical benefits for intelligent services and next-generation networks like 6G. Previous works have explored semantic compression techniques [1], knowledge-enhanced transmission strategies [2], and semantic-aware resource allocation. These efforts have shown the promise of semantic communication for improving efficiency and robustness across diverse intelligent systems and communication environments.
While these works largely emphasize semantic compression and transmission optimization, our focus shifts to a distinct but complementary problem: logical deduction of hypotheses under resource constraints. In many real-world scenarios—such as distributed sensing or privacy-sensitive environments—the transmitter lacks knowledge of the receiver’s precise goal. Consequently, the challenge becomes transmitting the most semantically content-informative message possible. This framework has the potential to empower a variety of applications involving logical deduction: drones in search-and-rescue operations can share logical summaries of their local observations [3]; wearable health monitors can transmit structured symptom patterns to support remote diagnostics [4]; autonomous vehicles can communicate compact causal cues like “reduced traction ahead” [4]; and legal or educational systems can tailor logic-based explanations to user intent [5].
This paper builds on prior work that introduced a First-Order Logic (FOL)-based semantic communication framework for deductive inference. Drawing from the foundational ideas of Carnap [6] and Hintikka [7], we treated FOL as a formal language for representing and transmitting structured knowledge. While classical semantic systems were computationally intractable, our earlier work [8, 9, 10] addressed scalability by framing communication as a boolean-satisfiability model selection problem. Using model counting, we developed a bit-efficient, task-agnostic transmission strategy that enabled logical inference at the receiver, even when the transmitter lacked access to the downstream task.
In this paper, we take our efforts a step further and formalize semantic communication as a posterior distribution alignment problem over a logical space structured by First-Order Logic (FOL). The world is represented using a hierarchy of constituents, attributive constituents, and Q-sentences, and uncertainty is quantified via inductive logical probability over these structures. The transmitter, observing partial evidence about the true state, aims to select messages that maximize semantic informativeness—quantified via Carnap’s cont-information—without access to the receiver’s hypothesis set. We show that optimal message selection reduces to minimizing the KL divergence between transmitter and receiver posteriors over constituents, which in turn becomes a constrained resource allocation problem over observed attributive constituents. To make the analysis tractable, we apply Stirling’s approximation and exploit the symmetry of the resulting objective, deriving a globally optimal solution using a truncated water-filling algorithm. We extend this framework to multi-round communication, showing that repeated optimal allocations converge to an ideal posterior configuration under cumulative constraints. This formulation enables principled message selection under realistic communication budgets and, under the Bayesian MAP criterion at the receiver, minimizes error in hypothesis deduction. We derive the convergence rate for the multi-round setup and show that the algorithm converges linearly. Finally, we validate our theoretical findings on a custom deduction dataset, demonstrating superior performance in hypothesis selection compared to prior work, greedy selection, and random baselines.
The contribution of this paper is twofold. First, it introduces a formal connection between semantic communication and logical inference by framing message transmission as the problem of posterior distribution alignment between transmitter and receiver in a First-Order Logic framework. This perspective distinguishes the present work from prior approaches that emphasize semantic compression or task-specific heuristics, and provides a principled way to quantify semantic informativeness through Carnap’s cont-information and distributional divergence. Second, we develop and analyze a resource allocation formulation of this problem, showing that the optimal message composition follows a truncated water-filling solution. We establish that this strategy achieves global optimality in the relaxed setting and extend it to multi-round communication, where convergence properties are derived. Simulation results confirm that aligning posterior distributions yields significant improvements in hypothesis deduction accuracy compared to prior semantic communication strategies. Together, these contributions advance semantic communication beyond compression-oriented formulations and provide a theoretically grounded approach for logic-based deductive reasoning under communication constraints. In the next section, we introduce FOL representation of the world and problem setup.
II Related Work
Semantic and goal-oriented communication systems, inspired by Weaver’s Level B notion, focus on transmitting meaning rather than symbols. DeepSC and DeepJSCC [11, 12] pioneered autoencoder-based designs, while other works introduced task-oriented training [13, 14] and functional compression [15]. These methods achieve notable gains in efficiency but often rely on black training or static task definitions.
Beyond end-to-end learning, neuro-symbolic and graph-based approaches aim to combine logical reasoning with semantic transmission. Prior work [16, 17, 18, 19, 20] has shown that these methods strengthen semantic understanding by capturing structured meaning, making them promising foundations for communication systems. RuleTaker and ProofWriter [21, 22] demonstrated transformer-based reasoning, while recent studies integrate knowledge graphs and large models [23, 24] to encode structured semantics. Vector quantization [25] and generative models [26] further extend semantic communication by aligning semantic tokens with modulation schemes or reconstructing content via shared AI models.
Our work differs in that it does not rely on neural compression or task-specific encoders. Instead, we adopt a symbolic and probabilistic formalism grounded in First-Order Logic and inductive probabilities tailored to logical inference. This allows us to frame semantic communication for hypothesis deduction as a posterior distribution alignment problem, derive a globally optimal truncated water-filling allocation, and establish convergence guarantees. Thus, our contributions provide a theoretical complement to learning-driven and application-specific approaches in semantic communication.
III Background on FOL, Inductive Probabilities, and Semantic Information
III-A Representing the State of the World via First-Order Logic
We begin by illustrating how the state of the world can be represented using First-Order Logic (FOL). Consider an infinite world $\mathbf{W}_{I}$ characterized by a countably infinite set of entities $\mathbf{E}$ and a finite set of predicates $\mathbf{P}$. Each FOL sentence $m_{i}$ is described using the logical language $\mathcal{L}$, which includes the predicates $\mathbf{P}$, entities $\mathbf{E}$, quantifiers $\forall,\exists$, and logical connectives $\land,\lor,\neg$. To describe the state in $\mathbf{W}_{I}$, we start by introducing the concept of Q-sentences.
Definition 1 (Q-sentence).
A Q-sentence $Q_{i}(a,b)$ is defined as:
$Q_{i}(a,b)=\bigwedge_{r\in\mathcal{I}_{P}}\delta_{r}R_{r}(a,b)\land\delta^{\prime}_{r}R_{r}(b,a),$ (1)
where $a,b\in\mathbf{E}$ and $\delta_{r},\delta^{\prime}_{r}\in\{+1,-1\}$ indicate the presence or negation of predicates.
Building upon Q-sentences, we define attributive constituents to capture the relationships between individual entities.
Definition 2 (Attributive Constituent).
An attributive constituent $C_{z}(x)$ in $\mathbf{W}_{I}$ is:
$C_{z}(x)=\bigwedge_{i\in\mathcal{I}_{\mathcal{Q}}}\delta_{k}\{(Q_{i}(x,b)):\exists b\},$ (2)
where $\delta_{i}\in\{+1,-1\}$, $\mathcal{I}_{\mathcal{Q}}$ is the index set over Q-sentences, and $i=1,...|\mathcal{I}_{\mathcal{Q}}|$, and $b\in\mathbf{E}$. The expression encapsulates all relationships entity $x$ has within the world, expressing what kind of individual $x$ is.
To represent the state in an infinite world (in a generalized manner), we define constituents as combinations of attributive constituents.
Definition 3 (Constituent).
A constituent $C^{w}$ of width $w$ in $\mathbf{W}_{I}$ is a conjunction of attributive constituents:
$C^{w}=\bigwedge_{t=1}^{w}\;\{C_{z}(x):\exists x\},$ (3)
which represents the relational structure of entities in the world $\mathbf{W}_{I}$. Here, the width $w$ is the number of attributive constituents $C_{z}(x)$ present in the conjunction forming $C^{w}$.
Theorem 1 (FOL Representation in Infinite World [7]).
Any FOL statement $m$ in $\mathbf{W}_{I}$ can be represented as:
[…]
where $\mathcal{I}_{\mathcal{C}}$ is the index set of constituents.
In FOL-based representation of world $\mathbf{W}_{I}$, we introduce a probability measure $\mathcal{P}_{I}$ over the fundamental elements, i.e., constituents $C^{j}$. To define this measure, we make use of inductive logical (a.k.a., epistemic) probabilities, which serve as inductive Bayesian posteriors, built upon observations and a prior distribution if any.
Definition 4 (Inductive Logical Probability [6]).
For any statement $g$ and evidence $e$ (observations) in $\mathbf{W}_{I}$, the inductive logical probability (a.k.a., degree of confirmation) $c(g,e)$ that evidence $e$ supports statement $g$ is defined as:
$c(g,e)=p(g\mid e)=\frac{p(g\land e)}{p(e)}=\frac{l_{g}+\lambda(w)/w}{l+\lambda(w)},$ (5)
where $l$ is the number of observations, $l_{g}$ is the number of observations confirming statement $g$, $w$ is the width of the constituent $C^{w}$ instantiated by statement $g$, $\lambda(w)\geq 0$ is the prior coefficient, possibly a function of $w$.
Then, the inductive logical probability measure $\mathcal{P}_{I}$ over $\mathcal{C}$ is defined as $\mathcal{P}_{I}=\{c(C^{j},e)\mid j\in\mathcal{I}_{\mathcal{C}}\}$ where $\sum_{j\in\mathcal{I}_{\mathcal{C}}}c(C^{j},e)=1$. To quantify the remaining uncertainty after new observations, Carnap proposed the concept of semantic ”cont(ent)-information”, defined as:
Definition 5.
The cont-information in a world $\mathbf{W}_{I}$ quantifies the informativeness of an observation $e$ to a specific observer by measuring the reduction in uncertainty about the SotW. Given a state $C^{j}$ and evidence $e$ (observations) in $\mathbf{W}_{I}$, cont-information is defined as:
$\text{cont}(C^{j};e)=1-c(C^{j},e)=1-p(C^{j}|e),$ (6)
In simple terms $\text{cont}(C^{j};e)$ measures the uncertainty reduction about constituent $C^{j}$ being the true SotW after observing e. Next, the cont-information measure is used to devise a semantic communication framework.
IV Node-to-Node Semantic Communication and Deductive Inference of Hypothesis
We consider a node-to-node semantic communication setting in which a transmitter node $N_{Tx}$ communicates with a receiver node $N_{Rx}$. The receiver’s objective is to identify the most plausible hypothesis $h^{*}\in\mathcal{H}=\{h_{i}\mid i\in\mathcal{I}_{\mathcal{H}}\}$, where each hypothesis $h_{i}$ is a self-consistent FOL statement expressible as in (4) and $\mathcal{I}_{\mathcal{H}}$ is the index set over hypotheses. The true world state corresponds to an unknown constituent $C^{*}\in\mathcal{C}$, and the receiver infers $h^{*}$, the hypothesis most consistent with $C^{*}$ based on messages $\{g_{t}\}_{t=1}^{T}$ received from the transmitter over $T$ rounds.
Previously, we introduced SCLD [8], a cont-information based message selection framework. At each round $t$, the transmitter, using its local observations $\mathcal{E}_{t}$, selects the semantically most content-informative FOL message $g_{t}$ of bounded size $|g_{t}|\leq B$, aiming to reduce the receiver’s uncertainty over $\mathcal{C}$ for better deduction of $h^{*}$. The transmitter does not know anything about the hypothesis set $\mathcal{H}$. As such, in our framework, the transmitter constructs and sends the message $g_{t}$ at round $t$ as:
$g_{t}^{*}=\mathop{\mathrm{arg\,max}}_{g\in\mathcal{E}_{t}\setminus\{g_{1},\dots,g_{t-1}\}}\text{cont}\left(g,\mathcal{E}_{r}\land\bigcup_{z=1}^{t-1}g_{z}\right),\text{s.t. }|g|\leq B,$ (7)
where $\mathcal{E}_{t}$ is the evidence observed by the transmitter and $\mathcal{E}_{r}$ is the evidence present at the receiver about the SotW.
Another approach to semantic communication for hypothesis inference involves accurately conveying the underlying distribution through messages $g$. In other words, the receiver’s objective becomes replicating the sender’s intended distribution as precisely as possible. This accuracy requirement can be quantitatively defined using the Kullback–Leibler (KL) divergence between the posterior distributions at the transmitter, $p(C^{w}_{t}\mid\mathcal{E}_{t})$, and at the receiver, $p(C^{w}_{r}\mid\mathcal{E}_{r})$:
$\mathrm{D}_{\mathrm{KL}}\!\left(p_{t}\,\|\,p_{r}\right)=\sum_{w=f}^{K}p_{t}(w)\log\frac{p_{t}(w)}{p_{r}(w)}.$ (8)
The transmitter then selects $g$ to minimize this divergence.
Upon receiving $g_{t}^{*}$ at round $t$, the receiver updates its inductive logical probability distribution $\mathcal{P}_{I}$ over constituents, refining its epistemic beliefs. After $T$ rounds, the receiver selects the most probable hypothesis as:
$h^{*}=\mathop{\mathrm{arg\,max}}_{h\in\mathcal{H}}c\left(h,\bigcup_{t=1}^{T}g_{t}\right).$ (9)
To analyze our framework, we consider finite evidence $\mathcal{E}$ containing $f$ distinct attributive constituents about the SotW, and let $n_{j}$ be the number of attributive constituents corresponding to $C_{j}(x)$, i.e., the number of times $n_{j}$ that has shown up in set $\mathcal{E}$. A constituent $C^{w}$ of width $w$ is compatible with $\mathcal{E}$ if $f\leq w\leq K$, where $f$ is the width of the minimal constituent and $K=4^{|\mathbf{P}|\times|\mathbf{E}_{\mathcal{E}}|^{2}}/2$ is the total number of possible attributive constituents, where $\mathbf{E}_{\mathcal{E}}$ is the set of unique entities in $\mathcal{E}$. The posterior over $\mathcal{C}$, the set of constituents, is given by:
$\mathcal{P}_{I}=\{c(C^{w},\mathcal{E})\mid f\leq w\leq K\},$ (10)
with each term derived via Bayes’ rule:
$p(C^{w}\mid\mathcal{E})=\frac{p(C^{w})p(\mathcal{E}\mid C^{w})}{\sum_{w^{\prime}}p(C^{w^{\prime}})p(\mathcal{E}\mid C^{w^{\prime}})},$ (11)
If the prior $p(C^{w})$ and likelihood $p(\mathcal{E}\mid C^{w})$ are selected as in [7], the posterior $p(C^{w}\mid\mathcal{E})$ becomes:
$p(C^{w}\mid\mathcal{E})=$ (12)
where, $\alpha\geq 0$ is the inductive generalization coefficient. $\pi(\alpha,x)$ is a Pochhammer symbol (shifted factorial) defined as:
$\pi(\alpha,x)=\prod_{i=0}^{\alpha-1}(i+x)=\frac{\Gamma(x+\alpha)}{\Gamma(x)}.$ (13)
Here, $\Gamma(x)$ is the Gamma function. The expression in (12) involves ratios of these Pochhammer symbols (and thus Gamma functions), making it complex.
However, this form is amenable to asymptotic analysis using Stirling’s approximation, which describes the behavior of the Gamma function for large arguments: $\Gamma(z)\approx\sqrt{2\pi}\,z^{z-1/2}e^{-z}$ for $z\gg 0$. This approximation is generally accurate for $z\geq 1$ and very close for $z\geq 10$. Assuming large $\alpha$ (i.e., $\alpha\geq 1$) and/or large observation counts $n_{j}$ (i.e., $n_{j}\geq 1$), we can apply Stirling’s approximation to the Gamma function terms within the Pochhammer symbols in (12).
After applying the approximation and focusing on the dominant terms (where many constants and exponential factors cancel out), the posterior $p(C^{w}\mid\mathcal{E})$ simplifies considerably. The key dependencies involving the counts $n_{j}$ become apparent in the simplified form:
$p(C^{w}\mid\mathcal{E})\approx$ (14)
where $\alpha\geq 1$ and $\lambda\geq 0$ are constant coefficients determining the impact of priors in singular inductive inferences and inductive generalizations [7]. Next, we formulate an optimization objective to determine an optimal method of semantic communication for hypothesis deduction.
V Single Round Semantic Communication as Distribution Approximation
In formal semantics grounded in first-order logic (FOL), world representations possess a well-defined structure. This structure, as declared in (1), (2), (3), and (4), permits any FOL statement including the hypothesis to have a well-defined probability assignment. Under this framework, the semantics of a hypothesis and it’s degree of confirmation can be determined. Hence, the optimal strategy for semantic communication for hypothesis deduction can be naturally formulated as the accurate transfer of this distribution. That is, the ultimate goal is for the receiver to reconstruct the intended distribution as accurately as possible. This fidelity criterion can be formalized using the Kullback–Leibler (KL) divergence between the transmitter $p_{t}=p(C^{w}_{t}\mid\mathcal{E}_{t})$ and receiver $p_{r}=p(C^{w}_{r}\mid\mathcal{E}_{r})$ posterior distributions:
$\mathcal{L}_{\text{KL}}=\mathrm{D}_{\mathrm{KL}}\left(p_{t}\,\big{\|}\,p_{r}\right)=\mathbb{C}(p_{t},p_{r})-\mathbb{H}(p_{t})$ (15)
where $\mathbb{H}(\cdot)$ is Shannon entropy and $\mathbb{C}(\cdot)$ is cross-entropy. Since $\mathbb{H}(p_{t})$ is constant w.r.t. the transmission, minimizing $\mathcal{L}_{\text{KL}}$ is equivalent to minimizing $\mathbb{C}(p_{t},p_{r})$, or maximizing:
$\max_{\{n_{j}^{\prime}\}}\sum_{w=f}^{K}p_{t}(w)\cdot p_{r}(w)\quad\text{s.t.}$ (16)
$\sum_{j=1}^{f}n_{j}^{\prime}=B,\quad 0\leq n_{j}^{\prime}\leq n_{j},\quad n_{j}^{\prime}\in\mathbb{Z}_{\geq 0}$ (17)
Here, $\{n_{j}^{\prime}\}$ are the counts of the $f$ distinct attributive constituent types, as determined by the selection of message $g$, to be transmitted to the receiver (forming $\mathcal{E}_{r}$), $B$ is the total transmission budget (e.g., total number of attributive constituents communicated), and $n_{j}$ is the available number of attributive constituent $C_{j}(x)$ at the transmitter (from $\mathcal{E}_{t}$). The transmitter distribution $p_{t}(w)=p(C^{w}_{t}\mid\mathcal{E}_{t})$ is fixed throughout communication as no new evidence is observed. The receiver distribution $p_{r}(w)=p(C^{w}_{r}\mid\mathcal{E}_{r})$ depends on the chosen $\{n_{j}^{\prime}\}$ via the approximate posterior relation derived from (14):
$p_{r}(w)\propto\left[\sum_{i=0}^{K-f}A_{w}(i)\cdot\prod_{j=1}^{f}(n_{j}^{\prime})^{\beta_{i}(w)}\right]^{-1}$ (18)
with $\beta_{i}(w):=\left(\frac{1}{f+i}-\frac{1}{w}\right)\lambda$. Notably, $\beta_{i}(w)$ is independent of the constituent type index $j$, making the objective function (16) symmetric in the variables $\{n_{j}^{\prime}\}$.
This is a constrained resource allocation problem (of distributing the limited transmission budget $B$ across possible attributive constituents to best approximate the transmitter’s belief distribution) under integer constraints, typical in communication systems where discrete packets or symbols must be transmitted. To find the optimal allocation, we analyze its continuous relaxation (treating $n_{j}^{\prime}$ as real, temporarily ignoring $n_{j}^{\prime}\leq n_{j}$). The Lagrangian is:
$\mathcal{L}(\{n_{j}^{\prime}\},\mu)$ $=\sum_{w=f}^{K}p_{t}(w)\cdot p_{r}(w)-\mu\left(\sum_{j=1}^{f}n_{j}^{\prime}-B\right)$ (19)
The Karush-Kuhn-Tucker (KKT) conditions require $\frac{\partial\mathcal{L}}{\partial n_{k}^{\prime}}=0$ for all $k=1,\dots,f$. Computing the derivative yields a system of equations relating $\{n_{j}^{\prime}\}$ and the Lagrange multiplier $\mu$. Crucially, due to the symmetry of the objective function (18) and the budget constraint $\sum n_{j}^{\prime}=B$, the expressions for $\frac{\partial\mathcal{L}}{\partial n_{k}^{\prime}}$ are identical for all $k$. Therefore, the unique solution satisfying all KKT conditions simultaneously must be symmetric:
$n_{j}^{\prime*}=\frac{B}{f},\quad\text{for all }j=1,\dots,f$ (20)
Given the symmetry, this stationary point corresponds to the global optimum of the relaxed problem [27]. This suggests evenly distributing the transmission budget $B$ across all $f$ relevant semantic constituent types when capacities are unlimited.
In practice, we must respect the integer nature of counts and the per-attributive-constituent type capacity limits $n_{j}^{\prime}\leq n_{j}$, and hence (19) cannot be directly implemented. Instead, incorporating these constraints transforms the solution into a truncated water-filling allocation. To find the optimal integer allocation $\{n_{j}^{\prime*}\}$ satisfying (17), the constituent types are first sorted based on available counts ($n_{(1)}\leq\dots\leq n_{(f)}$). Next, a water level $\theta$ is determined such that $\sum_{j=1}^{f}\min(\theta,n_{j})=B$, yielding the optimal continuous allocation respecting capacities, $n_{j}^{\text{cont}}=\min(\theta,n_{j})$. Finally, the integer allocation $\{n_{j}^{\prime*}\}$ is obtained by rounding $\{n_{j}^{\text{cont}}\}$ (e.g., using a standard rounding procedure) such that $\sum_{j=1}^{f}n_{j}^{\prime*}=B$ and $n_{j}^{\prime*}\leq n_{j}$ for all $j$. This algorithm provides the optimal message composition strategy for single-round semantic communication under budget and evidence capacity constraints.
Lemma 1 (Rounding Error Bound).
Let $\{n_{j}^{\mathrm{cont}}\}$ denote the optimal continuous allocation obtained by truncated water-filling and $\{n_{j}^{\prime*}\}$ the corresponding integer allocation after rounding subject to $\sum_{j}n_{j}^{\prime*}=B$. Then the deviation satisfies
$\max_{j}\left|n_{j}^{\prime*}-n_{j}^{\mathrm{cont}}\right|\leq 1,$ (21)
and consequently the cumulative rounding error is bounded as
$\left\|\mathbf{n}^{\prime*}-\mathbf{n}^{\mathrm{cont}}\right\|_{1}\leq f.$ (22)
Proof.
Since rounding each component to the nearest integer introduces an error of at most $1$, and the budget constraint ensures that positive and negative deviations must balance across $f$ components, the $\ell_{\infty}$ bound follows directly, with the $\ell_{1}$ bound obtained by summation. ∎
Remark 1 (Symmetry Assumption).
The global optimality of the symmetric allocation in (20) follows under the assumption that the feasible set is permutation invariant, i.e., the upper bounds on $n_{j}^{\prime}$ are identical across all $j$. When heterogeneous bounds $n_{j}$ are present, the continuous stationary point remains symmetric but may no longer lie in the feasible set. In such cases, the optimal allocation is obtained via truncated water-filling, which respects individual capacity limits while preserving the symmetry argument within the feasible region.
VI Multi-Round Semantic Communication
In many practical systems, communication occurs over multiple rounds. We extend our single-round water-filling framework to a multi-round setting with $T$ rounds, where each round $t$ has a transmission budget $B$. The optimal strategy depends on when the receiver performs hypothesis deduction. We consider two primary scenarios:
I. Long-term Optimization (Final Hypothesis Selection): If the receiver evaluates hypotheses only after all $T$ rounds are completed, the transmitter’s goal is to optimize the quality of the final, cumulative message set. The total budget available across all rounds is $TB$. This scenario effectively reduces to the single-round optimization problem (Section V) but with an increased total budget. The optimal cumulative allocation $\{n_{j}^{\prime\text{total}}\}$ is found using the truncated water-filling algorithm with total budget $TB$ and the initial available capacities $\{n_{j}\}$. Specifically, find the water level $\theta_{\text{total}}$ such that $\sum_{j=1}^{f}\min(\theta_{\text{total}},n_{j})=TB$, and the corresponding allocation is $n_{j}^{\prime\text{total}}=\min(\theta_{\text{total}},n_{j})$ (followed by integer rounding respecting constraints).
II. Short-term Optimization (Per-Round Hypothesis Selection): If the receiver performs hypothesis deduction after each round $t$, the transmitter must optimize the message sent in the current round to maximize the receiver’s inference quality incrementally, considering the information already transmitted. At each round $t$ ($1\leq t\leq T$), the transmitter calculates the optimal allocation $\{n_{j}^{\prime(t)}\}$ for the current round using the truncated water-filling algorithm with the per-round budget $B$ and the remaining capacities $\{r_{j}(t)\}$. The remaining capacity for constituent type $j$ at the start of round $t$ is:
$r_{j}(t)=n_{j}-\sum_{\tau=1}^{t-1}n_{j}^{\prime(\tau)}$ (23)
where $n_{j}^{\prime(\tau)}$ was the allocation for type $j$ transmitted in round $\tau$. The water level $\theta_{t}$ for round $t$ satisfies $\sum_{j=1}^{f}\min(\theta_{t},r_{j}(t))=B$, and the allocation for round $t$ is $n_{j}^{\prime(t)}=\min(\theta_{t},r_{j}(t))$ (followed by integer rounding respecting the budget $B$ and remaining capacities $r_{j}(t)$). This myopic strategy maximizes the immediate semantic gain at each round under the per-round budget constraint.
VII Message Selection Algorithm from Candidates
The strategies above assume the transmitter can construct the exact water-filling allocation. In practice, the transmitter might have a predefined set of $M$ candidate messages, $\mathcal{M}=\{\mathcal{A}^{(1)},\dots,\mathcal{A}^{(M)}\}$. For example, these message can be identical to the M evidences that the transmitter has observed. Each candidate message $\mathcal{A}^{(m)}=\{n_{j}^{(m)}\}_{j=1}^{f}$ represents a valid allocation of semantic constituents for a single round, satisfying $\sum_{j=1}^{f}n_{j}^{(m)}=B$ and $n_{j}^{(m)}\leq n_{j}$ (assuming capacities $n_{j}$ are known or estimated). The transmitter must select $T^{\prime}$ ($T^{\prime}\leq T$) messages from $\mathcal{M}$ for transmission over $T^{\prime}$ rounds. This selection should approximate the ideal strategies outlined in Section VI.
Again, the selection algorithm depends on the optimization objective:
a) Approximate Long-Term Optimization: To approximate the ideal long-term strategy, the transmitter selects a subset $\mathcal{S}\subset\mathcal{M}$ of size $|\mathcal{S}|=T^{\prime}$ such that the final cumulative allocation $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ is as close as possible to the ideal total water-filling allocation $\mathrm{WF}^{\text{final}}$ calculated with budget $T^{\prime}B$. The objective is to find the subset $\mathcal{S}$ that minimizes:
$\left\|\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}-\mathrm{WF}^{\text{final}}\right\|$ (24)
subject to the constraint that the cumulative allocation respects capacities: $\sum_{m\in\mathcal{S}}n_{j}^{(m)}\leq n_{j}$ for all $j$. In this case, the order of message transmission within the $T^{\prime}$ rounds does not affect the final outcome. This is a subset selection problem, which can be complex but aims to match the overall target distribution optimally using the available message pool.
b) Approximate Short-Term Optimization: To approximate the ideal per-round optimization strategy, the transmitter should select messages greedily over $T^{\prime}$ rounds. At each round $t$ ($1\leq t\leq T^{\prime}$), let the cumulative allocation sent so far be $\mathbf{N}^{(t-1)}=\sum_{\tau=1}^{t-1}\mathcal{A}^{(\sigma(\tau))}$, where $\sigma(\tau)$ is the index of the message sent in round $\tau$. The ideal target cumulative allocation after $t$ rounds, according to the ideal short-term strategy, is $\mathrm{WF}^{(t)}$, representing the water-filling allocation with total budget $tB$. The transmitter selects the next message $\mathcal{A}^{(\sigma(t))}$ from the remaining candidates in $\mathcal{M}$ (ensuring capacities are not exceeded, i.e., $N_{j}^{(t-1)}+n_{j}^{(\sigma(t))}\leq n_{j}$) that minimizes the deviation from the ideal target for the current cumulative step, e.g., minimizing $\|\mathbf{N}^{(t-1)}+\mathcal{A}^{(\sigma(t))}-\mathrm{WF}^{(t)}\|$. This involves selecting both the message and its transmission order $\sigma$. A common heuristic is to greedily choose the message at round $t$ that brings the current cumulative allocation $\mathbf{N}^{(t)}$ closest to the ideal target $\mathrm{WF}^{(t)}$.
VIII Convergence Analysis
We now analyze the convergence properties of the practical message selection algorithms presented in Section VII, comparing their performance to the ideal water-filling allocations derived in Section VI. The analysis considers the two optimization objectives: minimizing cumulative error over rounds (short-term) and minimizing final error (long-term). Let $T^{\prime}$ be the number of messages selected and transmitted.
Lemma 2 (Convergence Rate for Short-Term Optimization Approximation).
Let $\mathbf{N}^{(T^{\prime})}=\sum_{t=1}^{T^{\prime}}\mathcal{A}^{(\sigma(t))}$ denote the cumulative allocation after $T^{\prime}$ rounds using the greedy message selection algorithm (Section VII.a) which aims to approximate the ideal short-term strategy. Let $\mathrm{WF}^{(T^{\prime})}$ be the ideal cumulative water-filling allocation targeting budget $T^{\prime}B$. The maximum deviation between the achieved allocation and the ideal target satisfies:
$\max_{j=1,\dots,f}\left|N_{j}^{(T^{\prime})}-\mathrm{WF}_{j}^{(T^{\prime})}\right|=\mathcal{O}\left(\frac{f}{T^{\prime}}\right).$ (25)
Proof Sketch. The greedy algorithm at each round $t$ selects a message $\mathcal{A}^{(\sigma(t))}$ to minimize the deviation of the current cumulative sum $\mathbf{N}^{(t)}$ from the ideal target $\mathrm{WF}^{(t)}$. By iteratively reducing the largest deviations across the $f$ constituent types using messages of bounded size ($B$), the overall maximum deviation after $T^{\prime}$ rounds tends to decrease proportionally to $1/T^{\prime}$. The factor $f$ accounts for balancing across all constituent types.
Lemma 3 (Convergence Rate for Long-Term Optimization Approximation).
Let $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ be the final cumulative allocation achieved by selecting a subset $\mathcal{S}$ of $T^{\prime}$ messages from the candidate pool $\mathcal{M}$ using the algorithm in Section VII.b, aiming to approximate the ideal long-term strategy. Let $\mathrm{WF}^{\text{final}}$ be the ideal water-filling allocation targeting budget $T^{\prime}B$. The maximum deviation satisfies:
$\max_{j=1,\dots,f}\left|N_{j}^{\text{final}}-\mathrm{WF}_{j}^{\text{final}}\right|=\mathcal{O}(1).$ (26)
Proof Sketch. This algorithm selects the best fixed subset $\mathcal{S}$ of size $T^{\prime}$ from the candidate pool $\mathcal{M}$ to match the final target $\mathrm{WF}^{\text{final}}$. The approximation error is primarily determined by the granularity of the candidate messages in $\mathcal{M}$ and how well any combination of $T^{\prime}$ messages can match the specific target profile. Since the optimization only considers the final state and relies on a potentially limited discrete set of candidate messages, the error bound does not necessarily decrease with $T^{\prime}$ and remains bounded by a constant factor reflecting the quality and diversity of the pool $\mathcal{M}$.
IX Simulation Results and Discussion
Figure 1: Average Cosine Distance between receiver and transmitter distributions over $T=20$ rounds. Lower is better.
We evaluate the proposed semantic communication strategies through simulations designed to measure both the alignment between transmitter and receiver distributions and the resulting accuracy in hypothesis deduction. All derivations and simulations assume an error-free logical channel, i.e., transmitted constituents are received without distortion. This assumption isolates the effect of semantic resource allocation from channel impairments. In practice, symbol errors and partial misdecoding are unavoidable [28], and their impact on the proposed allocation strategy remains an open question. Preliminary robustness checks under small error rates indicate graceful degradation, and a systematic analysis of noisy channels constitutes an important direction for future work.
The system consists of a transmitter that observes $N_{obs}$ attributive constituents drawn from a distribution over $K$ types, while the receiver begins with no prior evidence. Communication proceeds for $T$ rounds, each constrained by a budget $B$ constituents per message. Unless otherwise stated, parameters are set to $K\approx 1000$, $N_{obs}=750$, $T=20$, $B=5$, $\lambda=1.0$, and the hypothesis set size $K_{\mathrm{hypo}}=10$. For each run, one hypothesis is selected as the ground truth and the transmitter’s evidence is generated from its associated distribution. All results are averaged over $N_{\mathrm{sim}}=50$ independent runs.
Five strategies are compared. The Random-Free and Random-Chunk methods serve as baselines, selecting constituents either directly from the transmitter’s pool or from predefined evidence chunks. The SCLD approach follows the semantic cont-information framework by selecting messages that maximize immediate informativeness. Finally, the WF-Greedy and WF-Long strategies implement the proposed truncated water-filling allocation, the former optimizing myopically at each round and the latter computing a global plan over the total budget $TB$.
Performance is quantified using Cosine distance between the transmitter’s and receiver’s distributions. Although many similarity measures are available in the literature—including Kullback–Leibler divergence, Jensen–Shannon distance, Total Variation distance, and Hellinger distance—it is well established that these measures provide closely related trends when comparing probability distributions. Cosine distance therefore serves as a representative and computationally efficient proxy. Extending the evaluation to multiple metrics may nevertheless provide additional insights into the sensitivity of the algorithms, and constitutes an important direction for future work.
Figure 1 shows the convergence of the receiver’s distribution toward the transmitter’s distribution over $T=20$ rounds. Both water-filling strategies significantly outperform the baselines, with WF-Greedy providing faster initial convergence and WF-Long achieving the lowest final divergence. WF-Greedy and WF-Long both reach $98\%$, compared to $89\%$ for SCLD and below $80\%$ for the random baselines. These results confirm that aligning posterior distributions leads directly to improved hypothesis inference.
Experiments further reveal that the relative differences in performance among the strategies depend on and vary with the communication budget $B$ and the number of samples $N_{obs}$. In some regimes, WF-Greedy provides stronger advantages, while in others WF-Long dominates. There are certain cases where random provides close accuracy to WF-Greedy and even surpassing WF-Long. This variability suggests that sensitivity analysis across a broader range of budgets, sample sizes, and evidence imbalance is necessary to fully characterize algorithmic performance. Moreover, although the present evaluation is limited to synthetic Dirichlet–multinomial environments, extending experiments to real-world reasoning datasets involving noise, incomplete evidence, and large-scale logical corpora and incorporating varying budget constraints will provide a more realistic assessment of the proposed framework. Despite these limitations, the initial results are highly promising, demonstrating the effectiveness of posterior alignment as a principle for semantic communication in hypothesis deduction.
X Conclusion
This paper analyzed semantic communication for logic-based hypothesis deduction by framing message transmission as posterior distribution alignment between transmitter and receiver. We showed that the resulting constrained resource allocation problem admits a globally optimal truncated water-filling solution, and extended the framework to multi-round communication with established convergence properties. Simulations demonstrated that aligning posteriors significantly improves hypothesis inference, with both greedy and long-term strategies outperforming random and cont-information baselines under the MAP rule, while remaining robust to evidence imbalance and moderate noise. Although our study was limited to synthetic Dirichlet–multinomial environments and assumed noiseless communication, the promising results motivate future work on scalability to larger logical corpora, robustness under noisy or incomplete evidence, and integration with practical constraints in next-generation semantic communication.